### v1

@@ Abstract
We address the challenge of product configuration in the context of increasing customer demand for diverse and complex products. We propose a solution through a curated selection of product model benchmarks formulated in the Coom language, divided into three fragments of increasing complexity. Each fragment is accompanied by a corresponding bike model example, and additional scalable product models are included in the CoomSuite, along with relevant resources. We outline an ASP-based workflow for solving Coom-based configuration problems, highlighting its adaptability to different paradigms and alternative ASP solutions. The CoomSuite aims to provide a comprehensive, accessible, and representative set of examples that can serve as a common ground for stakeholders in the field of product configuration.
Under consideration for publication in Theory and Practice of Logic Programming (TPLP)
@@ 1 Introduction
Configuration (Felfernig et al., 2014), a longstanding challenge in AI, is gaining renewed attention as manufacturing industries grapple with the challenge of meeting customer demands for high variance and complexity, all while operating under a high level of automation. To achieve this, configuration processes must empower customers to tailor products to their specific needs within a predefined configuration space. This space is designed to seamlessly integrate with subsequent processes (pricing, quotation, manufacturing, resource planning, and delivery), enabling them to be executed in a highly automated manner. Despite the existence of various approaches to product configuration (Felfernig et al., 2014), building and utilizing complex product models for this purpose remains a challenge in industrial settings. This is due to the diversity of methods, representations, and configuration systems available. Furthermore, stakeholders from industry, research, and software development often struggle to align their understanding of the practical application task with the precise semantics of the modeling formalism.
@@
To address this challenge, we propose a curated selection of product model benchmarks that encapsulate the key challenges of product configuration. These benchmarks can serve as a common ground for fostering a meaningful exchange among all stakeholders involved. More precisely, we formulate our benchmarks in the industrial configuration language Coom (Baumeister et al., 2025b), [^1] Coom is developed by denkbares and used in numerous industrial applications. and distinguish three language fragments of increasing complexity. The basic one, CoomCore, is mainly about (discrete) attributes and resembles a simple constraint satisfaction problem. The extended one, Coom[p], adds partonomies and cardinalities. And the last considered Coom fragment, Coom[x] adds numeric variables and calculations on top of Coom[p]. Each such class is accompanied by an exemplary product model from the domain of bikes, more precisely, a KidsBike, CityBike, and TravelBike model in Coom. Extensions for two of these language fragments called Coom[p*] and Coom[x*], respectively, enable modeling with unbounded cardinalities, where the latter is exemplified by the product model of a CargoBike. Further scalable product models are part of the CoomSuite (Baumeister et al., 2025a), which is intended to serve as a workbench for experimentation with industrial-scale product configuration problems, and contains all benchmark files, Coom grammar definitions, and further resources related to Answer Set Programming (ASP Lifschitz, 2019).
The workflow for solving Coom-based configuration problems with ASP is outlined in Figure 1.
@@
We start by converting Coom specifications in the form of a model and, optionally, a user input into ASP facts. This conversion is guided by a Coom grammar, which is supplied through the CoomSuite. A specialized ASP visitor then processes the resulting parse tree, translating it into a corresponding set of facts. This initial set of facts serves as a direct serialization of the tree structure and is further refined through an ASP encoding tailored to the specific ASP solution being implemented. The refined fact format, in conjunction with its corresponding configuration encoding, comprehensively defines the configuration space. Ultimately, the answer sets representing valid configurations are transformed back into Coom solutions. Additionally, the CoomSuite offers the possibility to interactively solve configuration problems through a Coom-specific but application-independent user interface (UI)
@@
The remainder of our paper is organized as follows: Section 2 introduces the CoomCore language and demonstrates its application through the KidsBike example. Section 3 presents our ASP-based approach to product configuration, detailing the refined fact format and ASP encoding. Section 4 outlines various extensions to the CoomCore language which contain commonly used features in product modeling and how they are handled within our ASP framework. These include concepts such as partonomy, cardinalities, and numerics, as well as user input and unbounded cardinalities. Further, an alternative encoding for solver fclingo is presented. Finally, Section 5 describes the CoomSuite UI along with its ASP implementation through system clinguin and Section 6 demonstrates the practical value of the CoomSuite by evaluating alternative ASP solutions against benchmarks of varying complexity levels.
@@
Early on, a general ontology for representing product configuration problems was introduced in (Soininen et al., 1998), which inspired many of Coom’s concepts. Coom also draws inspiration from the field of product ontologies (Lin et al., 1997; Baumeister, 2018). A wide range of approaches exist for representing and solving configuration problems across various paradigms (Junker, 2006; Hotz et al., 2014). In recent years, ASP has emerged as a promising alternative, as evidenced by several applications (Gebser et al., 2011; Felfernig et al., 2017; Gençay et al., 2019; Herud et al., 2022). Moreover, Falkner et al. (2015) developed an object-oriented approach to configuration by directly defining concepts in ASP, and Rühling et al. (2023) recently proposed an ASP-based approach rooted in firm mathematical foundations. In the context of interactive configuration, Falkner et al. (2020) conducted a comparative evaluation of various systems, including the ASP solver clingo as well as SAT and CP systems, for their suitability in this context, finding clingo to be as capable as any other system.
Another system for interactive configuration has been developed by Carbonelle et al. (2023), which uses the IDP-Z3 reasoning engine to interactively solve configuration problems specified in the FO()($\cdot$)( ⋅ ) language. This has been successfully applied to various applications such as finding suitable adhesives (Vandevelde et al., 2024) and machine component design (Aerts et al., 2022).
@@ 2 The CoomCore language
In this section, we introduce the CoomCore language fragment through the illustrative KidsBike example. This fragment primarily corresponds to basic constraint satisfaction problems (Dechter, 2003), focusing on variables that need to satisfy a given set of constraints. The KidsBike example, presented in Listing 1, encapsulates a simplified product configuration problem and incorporates all the essential language features of CoomCore.
@@ 14 W18 = ( 18 70 ) W20 = ( 20 80 )
Line 1 establishes the definition of a product. This represents the fundamental entity of the object we intend to configure. Note that only one product can be defined within a single Coom file. Inside a product several features are declared (Lines 2-5). A feature is declared by stating its type followed by its name, as in ‘Color color’. In this case, type and name are differentiated solely by capitalization. However, for the type Wheel, we have two distinct names: frontWheel and rearWheel. The type Bool is defined in Coom. In CoomCore, the only types of features are enumerations, representing finite domains of predefined choices. A simple example is the Color enumeration in Line 7. Furthermore, enumerations can include attributes to assign specific values to each option. For instance, the Wheel type defines two attributes: size and price (Line 10 and 11), and its options are declared in Lines 13-14. Unlike the Color enumeration, each Wheel option includes additional values for the size and price attributes, enclosed within parentheses. This indicates, for example, that the wheel W16 has a size of 16 and a price of 60.
Product features in CoomCore can be seen as variables and thus be used for expressing constraints between them. Coom also offers to use attributes of enumerations as variables, eg. by writing frontWheel.size for the size attribute of the feature frontWheel. In general, these variables are called path expressions.
One or more constraints can be declared within behavior structures. In Line 17, a conditional requirement is declared, stating that the size of the front wheel has to be larger than 16 when the color of the bike is Yellow. Such constraints are declared by using the keywords condition and require, followed by a simple logical (comparison) statement. These statements can contain path expressions and constants, numbers, or strings. In Line 23 a condition-free requirement is defined, stating that the size of the front and rear wheel have to be equal.
Another type of constraint is implemented through combination tables. The one in Lines 19-21 specifies that the use of a wheelSupport is obligatory with small wheels. More specifically, this constraint is declared using the keyword combinations in Line 19, followed by a list of path expressions that give the table’s column headers. Then, Lines 20 and 21 define the allowed combinations. A table entry can contain multiple values, as illustrated in the rearWheel column. Any combination that aligns with an allow line within the table represents a valid configuration for the product, while those that do not are invalid.
@@ 3 Solving Coom-based product configuration with ASP
Lines 7-9 give the (discrete) attributes of the configuration model. In fact, both the Wheel enumeration and one of its enumeration attributes size from the Coom model are represented as discrete attributes. The Bool attribute is a Coom built-in and added automatically during refinement. We omit showing the Color and price attribute. In Coom, the (combined) values of such attributes are among the enumeration’s options (cf. Section 2). We guarantee their compatibility, eg. of size and price from the Wheel enumeration, by means of a table constraint which is omitted for brevity.
Lines 11-22 give three constraints of the configuration model; they reflect Coom behavior[s]. Each constraint is identified with a term, c0, c1, etc., during refinement. Constraint c0 is a conditional requirement stating that the color Yellow is only available with larger wheels. For this, we require in Lines 11-14 that color=Yellow is false or frontWheel.size>16 is true. These statements are represented by f0 and f1, respectively. They are further decomposed via predicates binary/4 and unary/3 until the variable or constant level is reached. Analogously, Lines 21-22 encode a Boolean constraint stating that the size of the front and rear wheel must be equal.
@@
Next, Lines 4-17 encode Boolean and table constraints. To start with, the integrity constraint in Line 4 makes sure that all original constraints are satisfied. Lines 6-13 specify satisfaction conditions for various types of binary and unary formulas making up a Boolean constraint, viz. disjunction (Lines 6 and 7), negation (Line 8), and comparison operators = and > (Lines 10-13).
Lines 15-17 give satisfaction conditions for table constraints. For this, we identify all rows that are not satisfied (Lines 15 and 16). A row is not satisfied whenever one of its entries is not satisfied. Note that as entries can contain multiple values, we employ clingo’s conditional literals to check for all of the individual values within the table entry. Then, a table constraint is satisfied if one of its rows is satisfied (Line 17).
Listing 4 shows how to solve the KidsBike example with the solve mode of the CoomSuite following the workflow from Figure 1. The option --output coom (or -o coom for short) converts the ASP output to a (more readable) Coom format.
@@ 4 Coom language extensions
In this section, we show how various advanced configuration features can be modelled and solved using ASP within our framework.
The first part of the section displays how more commonly used features like partonomy, cardinalities, and numeric calculations can be modeled within Coom[x], using our TravelBike example for illustration. We do not delve into Coom[p], as all its concepts are already encompassed within Coom[x].
@@
Similar to above, we first outline how these features are modeled within Coom[x] and then demonstrate their representation and resolution using ASP. For this part, we first present the fact format and then proceed to show two similar but alternative encodings: One for solver clingo and one for fclingo. For both, we showcase selected aspects and refer the reader to Baumeister et al. (2025a) for the complete encoding.
@@
The second feature are unbounded cardinalities and allows modeling for cases where the exact (or approximate) number of some object is not known beforehand. Here, we introduce extensions to two of our language fragments which are called Coom[p*] and Coom[x*], respectively. Further, Coom[x*] is illustrated by the CargoBike example, a modified version of the TravelBike, where the number of possible bags is unbounded. For both, we proceed in the same manner as before by first showcasing a Coom example and then showing how this is solved using ASP.
@@ 12 B20 = ( 20 ) B50 = ( 50 ) B100 = ( 100 )
Listing 5 gives a simplified Coom[x] representation of the TravelBike example. First, the features of the bike are defined, starting with the two numeric ones, totalVolume and requestedVolume in Lines 2 and 3. Both are marked as such by the keyword num at the beginning of the line, followed by their respective ranges. The second feature can be thought of as a user requirement that can be set at runtime (cf. Section 4.4). Next, Lines 4 and 5 define the carrier and frame, whose types are captured by structure[s] in Lines 7 and 8. Other than enumeration[s], structure[s] may have features on their own, which allows for building complex partonomies. Both the Carrier and the Frame structure have exactly one feature, bag. The expressions 0..3 and 0..2 give their respective cardinalities. In fact, each feature in Coom has a cardinality, but when omitted it defaults to 1. In all our Coom fragments until now, lower and upper bounds are mandatory, and we carry this requirement over to our fact format as well (only in Section 4.5 we generalize this to allow for open bounds.) The type of both features is Bag, which in turn is an enumeration with a single attribute volume (Lines 10-13).
Lastly, we discuss the constraints of the model. The first constraint requires that there are no more than four bags in a configuration (Line 15). For this, it uses aggregate functions to count the number of bags. In general, aggregate functions perform calculations over a set of variables, which are defined implicitly in terms of a path expression, eg. carrier.bag. In this case, the function count(carrier.bag) returns the actual number of bags attached to the carrier in the configuration at hand.

### v2

@@ Abstract
We address the challenge of product configuration in the context of increasing customer demand for diverse and complex products. We propose a solution through a curated selection of product model benchmarks formulated in the Coom language, divided into three fragments of increasing complexity. Each fragment is accompanied by a corresponding example on bike configuration, and additional scalable product models are included in the CoomSuite, along with relevant resources. We outline an ASP-based workflow for solving Coom-based configuration problems, highlighting its adaptability to different paradigms and alternative ASP solutions. The CoomSuite aims to provide a comprehensive, accessible, and representative set of examples that can serve as a common ground for stakeholders in the field of product configuration.
Under consideration in Theory and Practice of Logic Programming (TPLP)
@@ 1 Introduction
Configuration (Felfernig et al. 2014), a longstanding challenge in AI, is gaining renewed attention as manufacturing industries grapple with the challenge of meeting customer demands for high variance and complexity, all while operating under a high level of automation. To achieve this, configuration processes must empower customers to tailor products to their specific needs within a predefined configuration space. This space is designed to seamlessly integrate with subsequent processes (pricing, quotation, manufacturing, resource planning, and delivery), enabling them to be executed in a highly automated manner. Despite the existence of various approaches to product configuration (Felfernig et al. 2014), building and utilizing complex product models for this purpose remains a challenge in industrial settings due to the diversity of methods, representations, and configuration systems available. Furthermore, stakeholders from industry, research, and software development often struggle to align their understanding of the practical application task with the precise semantics of the modeling formalism.
@@
To address this challenge, we propose a curated selection of product model benchmarks that encapsulate the key challenges of product configuration. These benchmarks can serve as a common ground for fostering a meaningful exchange among all stakeholders involved. More precisely, we formulate our benchmarks in the industrial configuration language Coom (Baumeister et al. 2025b), [^2] Coom is developed by denkbares and used in numerous industrial applications. and distinguish three language fragments of increasing complexity. The basic one, CoomCore, is mainly about (discrete) attributes and resembles a simple constraint satisfaction problem. The extended one, Coom[p], adds partonomies and cardinalities. And the last considered Coom fragment, Coom[x] adds numeric variables and calculations on top of Coom[p]. Each such class is accompanied by an exemplary product model from the domain of bikes, more precisely, a KidsBike, CityBike, and TravelBike model in Coom. Extensions for two of these language fragments called Coom[p*] and Coom[x*], respectively, enable modeling with unbounded cardinalities, where the latter is exemplified by the product model of a CargoBike. These and further scalable product models are bundled together in the CoomSuite (Baumeister et al. 2025a), which is intended to serve as a workbench for experimentation with industrial-scale product configuration problems, and in addition to all benchmark files, contains Coom grammar definitions, and further resources related to Answer Set Programming (ASP; Lifschitz 2019).
The workflow for solving Coom-based configuration problems with ASP is outlined in Figure 1. We start by converting product models in the form of a Coom specification and, optionally, a user input into ASP facts. This conversion is guided by a Coom grammar, which is supplied through the CoomSuite. A specialized ASP visitor then processes the resulting parse tree, translating it into a corresponding set of facts. This initial set of facts serves as a direct serialization of the tree structure and is further refined through an ASP encoding tailored to the specific ASP solution being implemented. The refined fact format, in conjunction with its corresponding configuration encoding, comprehensively defines the configuration space. Ultimately, the answer sets representing valid configurations are transformed back into Coom solutions.
@@
Additionally, the CoomSuite offers the possibility to interactively solve configuration problems through a Coom-specific but application-independent user interface (UI).
@@
The remainder of our paper is organized as follows: Section 2 introduces the CoomCore language and demonstrates its application through the KidsBike example. Section 3 presents our ASP-based approach to product configuration, detailing the refined fact format and ASP encoding. Section 4 outlines various extensions to the CoomCore language which contain commonly used features in product modeling and how they are handled within our ASP framework. These include concepts such as partonomy, cardinalities, and numerics, as well as user input and unbounded cardinalities. Further, an alternative encoding for solver flingo is presented. Finally, Section 5 describes the CoomSuite UI along with its ASP implementation through system clinguin and Section 6 demonstrates the practical value of the CoomSuite by evaluating alternative ASP solutions against benchmarks of varying complexity levels.
@@
Early on, a general ontology for representing product configuration problems was introduced in (Soininen et al. 1998), which inspired many of Coom’s concepts. Coom also draws inspiration from other ontologies aiming to represent corporate knowledge in a standardized way (Brunner et al. 2007; Coskun et al. 2009), especially in the context of products and industrial manufacturing (Lin et al. 1997; Baumeister 2018). A wide range of approaches exist for representing and solving configuration problems across various paradigms (Junker 2006; Hotz et al. 2014). In recent years, ASP has emerged as a promising alternative, as evidenced by several applications (Gebser et al. 2011; Felfernig et al. 2017; Gençay et al. 2019; Herud et al. 2022). Moreover, Falkner et al. (2015) developed an object-oriented approach to configuration by directly defining concepts in ASP, and Rühling et al. (2023) recently proposed an ASP-based approach rooted in firm mathematical foundations. In the context of interactive configuration, Falkner et al. (2020) conducted a comparative evaluation of various systems, including the ASP solver clingo as well as SAT and CP systems, for their suitability in this context, finding clingo to be as capable as any other system.
Another system for interactive configuration has been developed by Carbonnelle et al. (2023), which uses the IDP-Z3 reasoning engine to interactively solve configuration problems specified in the FO()($\cdot$) language. This has been successfully applied to various applications such as finding suitable adhesives (Vandevelde et al. 2024) and machine component design (Aerts et al. 2022).
Furthermore, in the related context of product-line variability modeling and configuration, the Clafer language (Juodisius et al. 2019) has been developed, which allows to capture in a unified way feature models (Felfernig et al. 2024), component models, and others. A series of supporting tools (Antkiewicz et al. 2013a) are available of which we highlight the interactive, web-based Clafer Configurator (Antkiewicz et al. 2013b).
@@ 2 The CoomCore language
In this section, we introduce the CoomCore language fragment through the illustrative KidsBike example. This fragment primarily corresponds to basic constraint satisfaction problems (Dechter 2003), focusing on variables that need to satisfy a given set of constraints. The KidsBike example, presented in Listing 1, encapsulates a simplified product configuration problem and incorporates all the essential language features of CoomCore.
@@ 14 W18 = ( 18 70 ) W20 = ( 20 80 )
Line 1 establishes the definition of a product. This represents the fundamental entity of the object we intend to configure. Note that only one product can be defined within a single Coom file. Inside a product several features are declared (Lines 2-5). A feature is declared by stating its type followed by its name, as in ‘Color color’. In this case, type and name are differentiated solely by capitalization. However, for the type Wheel, we have two distinct names: frontWheel and rearWheel. The type Bool is a built-in type in Coom. In CoomCore, a type of a feature must refer to an enumeration, representing finite domains of predefined choices. A simple example is the Color enumeration in Line 7. Furthermore, enumerations can include attributes to assign specific values to each option. For instance, the Wheel type defines two attributes: size and price (Line 10 and 11), and its options are declared in Lines 13-14. Unlike the Color enumeration, each Wheel option includes additional values for the size and price attributes, enclosed within parentheses. This indicates, for example, that wheel W16 has a size of 16 and a price of 60.
Features in CoomCore can be seen as variables and thus be used for expressing constraints over them. Coom also offers to use attributes of enumerations as variables, e.g., by writing frontWheel.size to refer to the size attribute of the feature frontWheel. In general, these variables are called path expressions.
Constraints can be declared within behavior structures. Here, we allow for three types of constraints. In Line 17, a conditional requirement is declared, stating that the size of the front wheel has to be greater than 16 if the color of the bike is Yellow. Such constraints are declared by using the keywords condition and require, followed by a simple logical (comparison) statement. These statements can contain path expressions and constants, numbers, or strings. In Line 23, a condition-free requirement is defined, stating that the size of the front and rear wheel have to be equal.
Another type of constraint is implemented through combination tables. In the example, Lines 19-21 specify that the use of a wheelSupport is obligatory with small wheels. More specifically, this constraint is declared using the keyword combinations in Line 19, followed by a list of path expressions as the table’s column headers. Then, Lines 20 and 21 define the allowed combinations. A table entry can contain multiple values, as illustrated in the rearWheel column. Any combination that aligns with an allow line within the table represents a valid configuration for the product, while those that do not are invalid.
@@ 3 Solving Coom-based product configuration with ASP
Lines 7-9 give the (discrete) attributes of the configuration model. In fact, both the Wheel enumeration and one of its enumeration attributes size from the Coom model are represented as discrete attributes. The Bool attribute is a Coom built-in and added automatically during refinement. We omit showing the Color and price attribute. In Coom, the (combined) values of such attributes are among the enumeration’s options (cf. Section 2). We guarantee their compatibility, e.g., of size and price from the Wheel enumeration, by means of a table constraint which is omitted for brevity.
Lines 11-22 give three constraints of the configuration model; they reflect Coom behavior[s]. Each constraint is identified with a term, c0, c1, etc., during refinement. Constraint c0 is a conditional requirement stating that the color Yellow is only available with larger wheels. For this, we require in Lines 11-14 that color=Yellow is false or frontWheel.size>16 is true. These statements are represented by f0 and f1, respectively, and the operators "||" and "!" stand for disjunction and negation, respectively. They are further decomposed via predicates binary/4 and unary/3 until the variable or constant level is reached. Analogously, Lines 21-22 encode a Boolean constraint stating that the size of the front and rear wheel must be equal.
@@
Next, Lines 4-17 encode Boolean and table constraints. To start with, the integrity constraint in Line 4 makes sure that all original constraints are satisfied. Lines 6-13 specify satisfaction conditions for various types of binary and unary formulas making up a Boolean constraint, viz. disjunction (Lines 6 and 7), negation (Line 8), and comparison operators = and > (Lines 10-13). To encode disjunction, we make use of clingo’s weight constraints, where the expression 1 <= { satisfied(XL); satisfied(XR) } states that at least one of the two subformulas XL or XR must be satisfied.
Lines 15-17 give satisfaction conditions for table constraints. For this, we identify all rows that are not satisfied (Lines 15 and 16). A row is not satisfied whenever one of its entries is not satisfied. Note that as entries can contain multiple values, we employ clingo’s conditional literals to check for all of the individual values within the table entry. Here, the expression not value(X,V) : allow(C,(Col,Row),V) will be expanded to not value(X,V1), not value(X,V2),... during grounding for a table entry containing the tuple (V1,V2,...). More precisely, the expression not value(X,V) : allow(C,(Col,Row),V) gathers all possible values V1,V2,... for a given table entry (Col,Row) of the table constraint C and then expands to a conjunction of literals not value(X,V1), not value(X,V2),.... Finally, a table constraint is satisfied if one of its rows is satisfied (Line 17).
Listing 4 shows how to solve the KidsBike example with the solve mode of the CoomSuite following the workflow from Figure 1. Currently, the CoomSuite uses clingo as the default solver. Here, kids-bike.coom is the input file containing the Coom model of the KidsBike example corresponding to Listing 1. The option --output coom (or -o coom for short) converts the ASP output to a (more readable) Coom format.
@@ 4 Coom language extensions
Having explained the basics of Coom, we now proceed to elaborate on its extensions and how they can be modelled and solved using ASP within our framework.
We begin by showing how more commonly used features like partonomy, cardinalities, and numeric calculations can be modeled within Coom[x], using our TravelBike example for illustration. We do not delve into Coom[p], as all its concepts are already encompassed within Coom[x].
@@
Similar to above, we first outline how these features are modeled within Coom[x] and then demonstrate their representation and resolution using ASP. For this part, we first present the fact format and then proceed to show two similar but alternative encodings: One for solver clingo and one for flingo. For both, we showcase selected aspects and refer the reader to Baumeister et al. (2025a) for the complete encoding.
@@
The second feature are unbounded cardinalities, which allow modeling cases where the exact (or approximate) number of some object is not known beforehand. Here, we introduce extensions to two of our language fragments, which we call Coom[p*] and Coom[x*], respectively. We illustrate Coom[x*] using the CargoBike example, a modified version of the TravelBike, where the number of possible bags is unbounded. For both, we proceed in the same manner as before by first showcasing a Coom example and then showing how this is solved using ASP.
@@ 12 B20 = ( 20 ) B50 = ( 50 ) B100 = ( 100 )
Listing 5 gives a simplified Coom[x] representation of the TravelBike example. First, the features of the bike are defined, starting with the two numeric ones, totalVolume and requestedVolume in Lines 2 and 3. Both are marked as such by the keyword num at the beginning of the line, followed by their respective ranges. The second feature can be thought of as a user requirement that can be set at runtime (cf. Section 4.4). Next, Lines 4 and 5 define the carrier and frame, whose types are captured by structure[s] in Lines 7 and 8. Both the num keyword and the structure keyword were not possible before in CoomCore. Other than enumeration[s], structure[s] may have features on their own, which allows for building complex partonomies. For example, the Carrier and the Frame structure have exactly one feature, bag. The expressions 0..3 and 0..2 give their respective cardinalities. Strictly speaking, each feature in Coom has a cardinality, but when omitted it defaults to 1..1. In all our Coom fragments until now, lower and upper bounds are mandatory, and we carry this requirement over to our fact format as well (only in Section 4.5 we generalize this to allow for open bounds.) The type of both features is Bag, which in turn is an enumeration with a single attribute volume (Lines 10-13).
Lastly, we discuss the constraints of the model. The first constraint requires that there are no more than four bags in a configuration (Line 15). For this, it uses aggregate functions to count the number of bags. In general, aggregate functions perform calculations over a set of variables, which are defined implicitly in terms of a path expression, e.g., carrier.bag. In this case, the function count(carrier.bag) returns the actual number of bags attached to the carrier in the configuration at hand.